OpenStudy (buttermequeasy):

What is the exact value of (sqrt) 38 / 2 (sqrt) 6? 19 (sqrt) 57/ 6 2 (sqrt) 57 / 3 (sqrt) 57 / 3 (sqrt) 57 / 6

OpenStudy (anonymous):

\[\sqrt{38} / (2\sqrt{6})\] is that your question?

OpenStudy (buttermequeasy):

Yes!

OpenStudy (anonymous):

do you know what a conjugate is?

OpenStudy (buttermequeasy):

No...?

OpenStudy (anonymous):

in easiest terms its a special fraction used to eliminate denominators of radicals and its always equivalent to unity (1)

OpenStudy (anonymous):

in your case the denominator is \[2\sqrt{6}\] so the conjugate is \[\sqrt{6} / \sqrt{6}\] which is equal to one, we together?

OpenStudy (buttermequeasy):

Yes.

OpenStudy (anonymous):

now multiply your question with the conjugate as follows which is the same as multiplying with one, so it'll not affect the value of your question but rearrange it as follows \[\left\{\sqrt{38} /2\sqrt{6} \right\} * \sqrt{6} /\sqrt{6}\] becomes:

OpenStudy (anonymous):

\[(\sqrt{38} * \sqrt{6} ) /2\]

OpenStudy (anonymous):

correction, it becomes \[(\sqrt{38} * \sqrt{6}) / 2*6\]

OpenStudy (anonymous):

we together butter...

OpenStudy (buttermequeasy):

Yes

OpenStudy (anonymous):

now multiply the two radicals 38 and 6 and you get \[(\sqrt{288} )/12\]

OpenStudy (anonymous):

now from your choices i can see root 57, implying that it must be a factor of 288 i.e 288 = 57 *4 , right?

OpenStudy (buttermequeasy):

Right.

OpenStudy (anonymous):

so \[\sqrt{288} = \sqrt{4} * \sqrt{57}\] but root 4 is equal to \[2\sqrt{57}\] are we toogether?

OpenStudy (anonymous):

so we've managed to reduce the numerator and we can now simplify the expression: \[(2\sqrt{57}) / 12\] divide 12 by 2 and get \[(\sqrt{57}) /6\]

OpenStudy (buttermequeasy):

Okay. Sorry this is a little confusing for me.

OpenStudy (anonymous):

where you lost?

OpenStudy (buttermequeasy):

The part where root 4 becomes root 2.

OpenStudy (anonymous):

 we together there?